# A Sampling-based Framework for Hypothesis Testing on Large Attributed Graphs [Scalable Data Science]

### v1

@@ 3.2.2. $\text{PHASE}_{\text{opt}}$ Algorithm
We further propose optimizations to reduce execution time. Algorithm 2 introduces two lines that replace line 7 in Algorithm 1. Firstly, in densely connected graphs, the computation of neighbor weights in step 8 of Algorithm 1 can be time-consuming due to the potentially large number of neighbors. To mitigate this, we randomly sample a subset of $\min\{|N|,n\}$ neighbors as candidates for selection, where $n$ is a parameter. Secondly, we adopt a non-backtracking approach by avoiding the selection of visited nodes during the sampling process(Lee et al., 2012). This strategy prevents cycles and unnecessary node visits.
@@
(showing only 2 lines that replace line 7 in Algorithm 1
1:$N^{\prime}\leftarrow N[v]-\mathcal{V_{S}}$ (Optim 2)
2:$N\leftarrow$ Select $\text{min}\{|N^{\prime}|,n\}\text{ nodes randomly from }N^{\prime}$ (Optim 1)
@@ 3.2.3. Convergence of Hypothesis Estimators
Let us demonstrate the convergence of hypothesis estimators using our proposed sampling algorithms (Challenge 1).
By Lemma 5.1 from(Ribeiro and Towsley, 2010), FrontierS is equivalent to the sampling process of a SRW over $\mathcal{G}^{m}=(\mathcal{V}^{m},\mathcal{E}_{m})$, where $\mathcal{G}^{m}$ is the $m$-th Cartesian power of $\mathcal{G}$ and
$\mathcal{V}^{m}=\{(v_{1},\ldots,v_{m})|v_{1}\in\mathcal{V}\wedge\dots\wedge v_{m}\in\mathcal{V}\}$
is the $m$-th Cartesian power of $\mathcal{V}$. $\forall\boldsymbol{u,v}\in\mathcal{V}^{m}$, $(\boldsymbol{u,v})\in\mathcal{E}_{m}$ if there exists $i$ such that $(u_{i},v_{i})\in\mathcal{E}$ and for all $j\neq i$, $u_{j}=v_{j}$. Given $\mathcal{G}$ is a directed, connected, and non-bipartite graph, in the steady state, the sampled edges of FrontierS form a stationary sequence and satisfy the following Strong Law of Large Numbers (SLLN), the proof of which can be found in (Ribeiro and Towsley, 2010).
@@
For a path hypothesis with length $l$, it involves $(l-1)$-th order random walk, which is equivalent to a SRW over $\mathcal{G}^{l-1}=(\mathcal{V}^{l-1},\mathcal{E}_{l-1})$(Lovász, 1993). Hence, PHASE’s random walks can be viewed as a SRW over $\mathcal{G}^{l-1+m}=(\mathcal{V}^{l-1+m},\mathcal{E}_{l-1+m})$. As PHASE ensures that the resulting random walks over $\mathcal{G}^{l-1+m}$ maintain both irreducibility and aperiodicity, it preserve stationarity and adhere to the SLLN. $\text{PHASE}_{\text{opt}}$ also follows the SLLN because the stationary distribution of the NBRW remains the same as that of the SRW.
By Theorem1, we can construct hypothesis estimators for node, edge, and path hypotheses and prove their convergence. Here, we focus on scenarios where agg represents an average function. The hypothesis estimators for other aggregation functions, such as maximum and minimum, can be derived in a similar manner.
A path hypothesis ($l\geq 2$): $\text{avg}(f_{\mathcal{P}}|M_{t_{i}},\forall t_{i}\text{ on }\mathcal{P})$ has a primary subject $f_{\mathcal{P}}$. Let $\mathcal{P}^{*}$ be all relevant paths in $\mathcal{G}$, and $\mathcal{P_{G}}$ be all paths with the same length as $\mathcal{P}$ in $\mathcal{G}$. The mean value of the path hypothesis, $\theta_{path}$, is
where $T(\mathcal{P})=f_{\mathcal{P}}\times\mathbbm{1}_{M_{t_{i}}\subseteq\mathcal{L}_{\phi(t_{i})}\forall t_{i}\text{ on }\mathcal{P}}$. Replacing $\mathcal{G}$ with the sampled graph $\mathcal{S}$, the estimator for $\theta_{path}$ is
When $B$ goes to infinity, $\mathcal{S}$ converges to $\mathcal{G}$. By Theorem1, we have
(3) $\lim_{B\to\infty}\frac{1}{|\mathcal{P}^{*}\cap\mathcal{P_{S}}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{S}}}T(\mathcal{P})\to\frac{1}{|\mathcal{P}^{*}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{G}}}T(\mathcal{P})$
almost surely. Hence, $\hat{\theta}_{path}$ is an asymptotically unbiased estimator of $\theta_{path}$. Hypothesis estimators for node and edge hypotheses can be derived similarly.
@@ 4. Experiments
The goal of our experiments is twofold: 1) test the effectiveness of optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE (Section 4.3), and 2) compare hypothesis-agnostic and hypothesis-aware samplers in terms of test significance (Section 4.4), accuracy (Section 4.5), and time (Section 4.6). The source code is publicly available  [^1] https://github.com/Carrieww/GraphHT.
@@ 4.1. Experimental Setup
MovieLens 9,705 996,656 1.06e-02
2
1
DBLP 1,623,013 11,040,170 4.19e-06
4
4
Yelp 2,136,118 6,743,879 1.48e-06
2
1
@@
Datasets. We used three datasets (Harper and Konstan, 2016; Tang et al., 2008), extracted from real attributed networks: MovieLens, DBLP, and Yelp. Table 1 shows the statistics of the datasets.
Samplers We compare $\text{PHASE}_{\text{opt}}$ with eleven existing samplers:
@@
  • Node samplers: Random Node Sampler (RNS), Degree-Based Sampler (DBS)
  • Edge samplers: Random Edge Sampler (RES)
@@
  • Random walk based samplers: Simple Random Walk (SRW), Frontier Sampler (FrontierS), Non-Backtracking Random Walk (NBRW), Random Walk with Restarter (RWR), Metropolis-Hastings Random Walk (MHRW), Snaw Ball Sampler (SBS), Forest Fire Sampler (FFS), Shortest Path Sampler (ShortestPathS)
Hypotheses. For each dataset (DB) and hypothesis type (N for node hypotheses), we provide three example hypotheses (labeled as DB-N1) in experiments. For example, on Yelp, we explore an edge hypothesis testing if “fast food average ratings exceed 4”. In DBLP, we examine a path hypothesis to see if “papers co-authored by Chinese institutes on data mining have over 50 citations”. To assess path length effects, we further test two path hypotheses with lengths of three and four in DBLP.
@@
Parameter Choice. For PHASE and $\text{PHASE}_{\text{opt}}$, We set $m=50$ to maintain the path-preserving capability of multi-dimensional random walks, and $n=30$ to strike a balance between accuracy and time efficiency. Additionally, we choose $w_{h}=10$ and $w_{l}=0.1$ to regulate the bias towards relevant nodes, edges, and paths. The sampling budget $B$ is maintained as a proportion of the total number of nodes in $\mathcal{G}$ for all samplers. As for existing hypothesis-agnostic samplers, we use the best parameters in their respective settings. We report an average of 30 runs for every evaluation measure.
@@ 4.2. Evaluation Measures
Accuracy. Accuracy at a specific sampling proportion reflects the effectiveness of a sampling method. The accuracy below measures the number of matched hypothesis testing results on $\mathcal{G}$ and $\mathcal{S}$:
@@
Time. We measure the total execution time, including both the sampling time and the time taken to extract relevant information from $\mathcal{S}$ for hypothesis testing.
Table 2. Average execution time (sec) of PHASE and $\text{PHASE}_{\text{opt}}$ for DBLP node, edge, and path hypotheses.
Node Edge Path
PHASE 115.66 539.15 441.3
$\text{PHASE}_{\text{opt}}$ 5.56 8.76 5.91
Table 3. The accuracy of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
Dataset
Hypothesis Type
Sampling Proportion (%)
$\text{PHASE}_{\text{opt}}$ RES RNS DBS SRW NBRW RWR MHRW ShortestPathS FrontierS FFS SBS
Node
1
0.9 0.89 0.83 0.86 0.88 0.89 0.87 0.89 0.86 0.84 0.83 0.56
Edge
2.5
0.98 0.68 0.77 0.98 0.99 1 0.99 0.98 0.99 0.73 0.93 0.91
MovieLens
Path
5
0.99 0.1 0.88 0.83 0.82 0.89 0.91 0.98 0.95 0.38 0.85 0.62
Node
0.2
0.96 0.48 0.87 0.93 0.92 0.91 0.9 0.94 0.92 0.92 0.94 0.88
Edge
0.2
0.76 0.48 0 0.71 0.73 0.69 0.7 0.29 0.69 0.63 0.7 0.42
DBLP
Path
0.2
0.89 0 0 0 0.26 0.3 0.32 0.18 0.33 0.043 0.3 0.12
Node
0.1
0.99 0.65 0.77 0.61 0.69 0.69 0.69 0.77 0.7 0.64 0.66 0.48
Edge
1
1 0.73 0.54 0.76 0.79 0.91 0.87 0.84 0.76 0.77 0.79 0.71
Yelp
Path
1
0.99 0.11 0.05 0.79 0.78 0.78 0.72 0.8 0.99 0.42 0.67 0.42
Table 4. The execution time (sec) of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
Dataset
Hypothesis Type
Sampling Proportion (%)
$\text{PHASE}_{\text{opt}}$ RES RNS DBS SRW NBRW RWR MHRWS ShortestPathS FrontierS FFS SBS
MovieLens
Node
1
0.083 0.99 0.023 0.077 0.057 0.06 0.05 0.06 0.063 0.083 0.067 0.047
Edge
2.5
0.45 0.99 0.11 0.36 0.31 0.36 0.33 0.29 0.26 0.33 0.35 0.31
Path
5
4.92 1.03 0.34 4.32 4.53 4.55 4.76 3.10 0.95 0.38 3.87 3.08
DBLP
Node
0.2
5.56 18.70 0.55 6.98 7.98 8.22 9.67 1.66 31.32 10.48 5.73 3.03
Edge
0.2
8.76 22.57 0.90 10.07 14.76 12.89 12.32 3.53 33.91 14.46 8.87 5.27
Path
0.2
5.44 19.61 0.71 6.87 8.24 8.39 8.13 2.17 31.13 9.37 5.33 3.01
Yelp
Node
0.1
2.56 13.84 1.02 6.81 1.19 0.96 1.16 1.12 1.84 1.68 1.02 6.58
Edge
1
19.42 16.88 2.01 13.61 8.58 8.97 9.17 6.93 30.23 10.22 10.27 6.03
Path
1
56.97 14.35 1.53 23.48 15.97 16.61 19.03 7.35 37.81 9.95 25.61 19.94

### v2

@@ 3.2.2. $\text{PHASE}_{\text{opt}}$ Algorithm
We further propose optimizations to reduce execution time based on heuristics. Algorithm 2 introduces two lines that replace line 7 in Algorithm 1. In line 1, we employ a non-backtracking approach to avoid selecting previously visited nodes during sampling (Lee et al., 2012), preventing cycles and unnecessary revisits. Additionally, in densely connected graphs, computing neighbor weights in line 8 of Algorithm 1 can be time-consuming due to the potentially large number of neighbors. To address this, line 2 in Algorithm 2 introduces random sampling of a subset of $\min\{|N^{\prime}|,n\}$ neighbors as candidates, where $n$ is a parameter. This can effectively reduce computation time. But it may result in some accuracy loss when $n$ is significantly smaller than $|N^{\prime}|$.
@@
(showing only 2 lines that replace line 7 in Algorithm  1
1:$N^{\prime}\leftarrow N[v]-\mathcal{V_{S}}$ (Optim 1)
2:$N\leftarrow$ Select $\text{min}\{|N^{\prime}|,n\}\text{ nodes randomly from }N^{\prime}$ (Optim 2)
@@ 3.2.3. Convergence of Hypothesis Estimators
We demonstrate the convergence of hypothesis estimators, which ensures the convergence of hypothesis testing accuracy to one, for all sampling methods. Then, we show our proposed sampler achieves earlier and smoother convergence (O2).
We will focus on scenarios where agg is an average function to construct and prove the convergence of hypothesis estimators. Estimators for other aggregation functions, such as maximum and minimum, can be derived analogously.
For a path hypothesis ($l\geq 0$), $\text{avg}(f_{\mathcal{P}}\mid M_{t_{i}},\forall t_{i}\text{ on }\mathcal{P})$ has a primary subject $f_{\mathcal{P}}$. Let $\mathcal{P}^{*}$ be all relevant paths in $\mathcal{G}$, and $\mathcal{P_{G}}$ be all paths with the same length as $\mathcal{P}$ in $\mathcal{G}$. The mean value of the path hypothesis, $\theta_{path}$, is
where $T(\mathcal{P})=f_{\mathcal{P}}\times\mathbbm{1}_{M_{t_{i}}\subseteq\mathcal{L}_{\phi(t_{i})}\forall t_{i}\text{ on }\mathcal{P}}$. Replacing $\mathcal{G}$ with the sampled graph $\mathcal{S}$, the estimator for $\theta_{path}$ on $\mathcal{S}$ is
When $l=0$ (resp. $l=1$), we name the corresponding mean value $\theta_{node}$ (resp. $\theta_{edge}$) and estimator $\hat{\theta}_{node}$ (resp. $\hat{\theta}_{edge}$).
Some sampling methods, including random edge sampler, simple random walk, and FrontierS, obey the Strong Law of Large Numbers (SLLN)  (Ribeiro and Towsley, 2010). Their $\hat{\theta}_{node}$ and $\hat{\theta}_{edge}$ are asymptotically unbiased estimators of $\theta_{node}$ and $\theta_{edge}$, respectively, when $B\to\infty$.
@@
In general scenarios with any path hypotheses ($l\geq 0$) and sampler, SLLN may not apply, and the estimator may not be asymptotically unbiased. However, $\hat{\theta}_{\text{path}}$ still converges to $\theta_{\text{path}}$ due to Finite Population Correction (FPC), a statistical adjustment made when sampling without replacement from a finite population (Bondy and Zlot, 1976). As $B$ increases, $\text{FPC}=\frac{|\mathcal{P}^{*}|-\mathcal{P_{S}}}{|\mathcal{P}^{*}|-1}$ approaches zero, which ensures the convergence of the hypothesis estimator to its ground truth.
While all samplers ensure the convergence of the hypothesis estimator, our proposed samplers achieve earlier convergence than other methods. This is due to their higher probability of selecting relevant nodes, edges, and paths. For instance, when $l=2$, the probability to pick nodes $(x_{1},x_{2},x_{3})$ is:
$P(x_{1},x_{2},x_{3})$ $=P(x_{1})\times(P(x_{2}\mid x_{1})\times A(x_{2}\mid x_{1}))$
×$\quad\times(P(x_{3}\mid x_{1},x_{2})\times A(x_{3}\mid x_{1},x_{2}))$
where $A(x_{i}\mid x_{i-1})$ indicates the accessibility (0 or 1) of node $x_{i}$ from $x_{i-1}$. Assuming node $x_{i}$ has $n_{i}$ neighbors with $k_{i}$ of type $x_{i+1}$ and there are $d$ $x_{1}$ in $m$ seed nodes, where $d\geq 0$ and $k_{i}\geq 0$:
$P_{\text{PHASE}}(x_{1},x_{2},x_{3})$ $=\frac{d\cdot w_{h}}{d\cdot w_{h}+(m-d)\cdot w_{l}}$
×$\quad\times\left(\frac{w_{h}}{k_{1}\cdot w_{h}+(n_{1}-k_{1})\cdot w_{l}}\times A(x_{2}\mid x_{1})\right)$
×$\quad\times\left(\frac{w_{h}}{k_{2}\cdot w_{h}+(n_{2}-k_{2})\cdot w_{l}}\times A(x_{3}\mid x_{1},x_{2})\right)$
Random node and edge samplers have the lowest probability of picking $(x_{1},x_{2},x_{3})$. Random walk based samplers have a transitional probability from $x_{i}$ of $\frac{1}{n_{i}}\leq\frac{w_{h}}{k_{i}\cdot w_{h}+(n_{i}-k_{i})\cdot w_{l}}$, particularly when $k_{i}$ is large. With a higher probability to pick $(x_{1},x_{2},x_{3})$, PHASE and $\text{PHASE}_{\text{opt}}$ have earlier convergence of the hypothesis estimator than existing hypothesis-agnostic samplers. Also, as shown in Figure 4, the empirical results of two path hypotheses, whose details are in Table 2, align with the claim.
Figure 4. The convergence of hypothesis estimator for two path hypotheses: DB-P3 (left) and YP-P3 (right).
@@ 4. Experiments
The goal of our experiments is twofold: 1) test the effectiveness of optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE (Section  4.3), and 2) compare hypothesis-agnostic and hypothesis-aware samplers in terms of test significance (Section  4.4), accuracy (Section  4.5), and time (Section  4.6).
@@ 4.1. Experimental Setup
MovieLens 9,705 996,656 1.06e-02 2 1
DBLP 1,623,013 11,040,170 4.19e-06 4 4
Yelp 2,136,118 6,743,879 1.48e-06 2 1
@@
Datasets. We use three datasets (Harper and Konstan, 2016; Tang et al., 2008; yel, 2015), extracted from real attributed networks: MovieLens, DBLP, and Yelp. Table 1 shows the statistics of the datasets.
Samplers We compare $\text{PHASE}_{\text{opt}}$ with 11 existing samplers:
@@
  • Node samplers: Random Node Sampler (RNS(Stumpf et al., 2005), Degree-Based Sampler (DBS)  (Adamic et al., 2001)
  • Edge samplers: Random Edge Sampler (RES)  (Krishnamurthy et al., 2005)
@@
  • Random walk based samplers: Simple Random Walk (SRW)  (Gjoka et al., 2010), Frontier Sampler (FrontierS)  (Ribeiro and Towsley, 2010), Non-Backtracking Random Walk (NBRW)  (Lee et al., 2012), Random Walk with Restarter (RWR)  (Leskovec and Faloutsos, 2006), Metropolis-Hastings Random Walk (MHRW)  (Stutzbach et al., 2009), Snaw Ball Sampler (SBS)  (Goodman, 1961), Forest Fire Sampler (FFS)  (Leskovec and Faloutsos, 2006), Shortest Path Sampler (ShortestPathS)  (Rezvanian and Meybodi, 2015)
Hypotheses. For each dataset (e.g., DB) and hypothesis type (e.g., N for node hypotheses), we chose three example hypotheses (e.g., DB-N1) in the experiment. Due to the page limit, examples for DBLP and Yelp are shown in Table 2. These hypotheses are chosen based on context and difficulty. Based on the context of the datasets, the subject of interest and the constant in the hypothesis should reflect meaningful requests from real users. The difficulty is related to the path length and the number of relevant nodes, edges, or paths, as indicated in the third column of Table 2. The longer the path or the fewer the relevant nodes, edges, or paths, the more difficult it becomes to sample them from $\mathcal{G}$ for accurate hypothesis testing. We use the following path hypotheses of DBLP with lengths of three and four in the experiment: “the average citation of two papers, each authored by a Microsoft researcher and citing each other, $>$ 50” and “the average citation of two conference papers, each authored by a Microsoft researcher, $>$ 50”.
Table 2. Examples of hypotheses for DBLP and Yelp
Hypothesis
Type
Example Relevant nodes,
edges, paths
DBLP
Node DB-N1: The avg citation of papers published as journals $>$ 20 199205 (easy)
DB-N2: The avg citation of papers in conferences in 2010 $>$ 10 31566 (medium)
DB-N3: The avg citation of papers published in Journal in 2017 $>$ 10 1588 (hard)
Edge DB-E1: The avg weight of conference papers on data mining $>$ 0.5 44925 (easy)
DB-E2: The avg weight of journal papers on data mining $>$ 0.5 13400 (medium)
DB-E3: The avg weight of papers on telecommunications network $>$ 0.5 2510 (hard)
Path DB-P1: The avg weight of papers by China’s institutes on data mining $>$ 0.5 17671 (easy)
DB-P2: The avg citation of papers co-authored by authors in Peking and China’s institutions $>$ 50 7065 (medium)
DB-P3: The avg citation of conference papers by Microsoft Researchers $>$ 50 3217 (hard)
Yelp
Node YP-N1: The avg reviews given by users with high popularity $>$ 200 112043 (easy)
YP-N2: The avg stars given by users who have low prolificacy and medium popularity $>$ 3 16429 (medium)
YP-N3: The avg number of reviews of business in Illinois $>$ 4 2144 (hard)
Edge YP-E1: The avg ratings of fast food $>$ 4 224536 (easy)
YP-E2: The avg ratings of furniture stores $>$ 3 33040 (medium)
YP-E3: The avg percentage of useful reviews given by useful writers to Illinois businesses $>$ 0.5 4242 (hard)
Path YP-P1: The avg rating difference on path [business in FL - high popularity user - business in LA] $>$ 0.5 615174 (easy)
YP-P2: The avg rating difference on path [business in LA - high popularity user - business in IL] $>$ 0.5 15542 (medium)
YP-P3: The avg rating difference on path [business in LA - medium popularity user - business in AB] $>$ 0.5 1080 (hard)
@@
Parameter Choice. We determine optimal settings for $m$, $n$, $w_{h}$, and $w_{l}$ through a grid search, with $m$ and $n$ ranging from 10 to 200, and $w_{h}$ and $w_{l}$ from 0.1 to 20. We set $m=50$ to maintain the path-preserving ability of multi-dimensional random walks, and $n=30$ to strike a balance between accuracy and time efficiency. $w_{h}=10$ and $w_{l}=0.1$ help regulate the prioritization towards sampling relevant nodes, edges, or paths. The sampling budget $B$ is maintained as a proportion of the total number of nodes in $\mathcal{G}$ for all samplers. In real deployment, we recommend iteratively increasing $B$ until the accuracy stabilizes at a high threshold (e.g., 0.9) based on the average from 30 samples to determine the optimal $B$. As for existing hypothesis-agnostic samplers, we use the best parameters in their respective settings. We report an average of 30 runs for every evaluation measure.
@@ 4.2. Evaluation Measures
Accuracy. Accuracy at a specific sampling proportion reflects the effectiveness of a sampling method. It measures the number of matched hypothesis testing results on $\mathcal{G}$ and $\mathcal{S}$:
@@
Time. We measure the total execution time, including the sampling time, the time taken to extract relevant information from $\mathcal{S}$ for hypothesis testing, and hypothesis testing time.
Table 3. The accuracy of 11 existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
Dataset Hypothesis Type Sampling Proportion (%) $\text{PHASE}_{\text{opt}}$ PHASE RES RNS DBS SRW NBRW RWR MHRW ShortestPathS FrontierS FFS SBS
Node 1 0.9 1 0.89 0.83 0.86 0.88 0.89 0.87 0.89 0.86 0.84 0.83 0.56
Edge 2.5 0.98 1 0.68 0.77 0.98 0.99 1 0.99 0.98 0.99 0.73 0.93 0.91
MovieLens Path 5 0.99 1 0.1 0.88 0.83 0.82 0.89 0.91 0.98 0.95 0.38 0.85 0.62
Node 0.2 0.96 1 0.48 0.87 0.93 0.92 0.91 0.9 0.94 0.92 0.92 0.94 0.88
Edge 0.2 0.76 0.79 0.48 0 0.71 0.73 0.69 0.7 0.29 0.69 0.63 0.7 0.42
DBLP Path 0.2 0.89 0.88 0 0 0 0.26 0.3 0.32 0.18 0.33 0.043 0.3 0.12
Node 0.1 0.99 1 0.65 0.77 0.61 0.69 0.69 0.69 0.77 0.7 0.64 0.66 0.48
Edge 1 1 1 0.73 0.54 0.76 0.79 0.91 0.87 0.84 0.76 0.77 0.79 0.71
Yelp Path 1 0.99 1 0.11 0.05 0.79 0.78 0.78 0.72 0.8 0.99 0.42 0.67 0.42
Table 4. The execution time (sec) of 11 existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
Dataset Hypothesis Type Sampling Proportion (%) $\text{PHASE}_{\text{opt}}$ PHASE RES RNS DBS SRW NBRW RWR MHRWS ShortestPathS FrontierS FFS SBS
MovieLens Node 1 0.083 0.41 0.99 0.023 0.077 0.057 0.06 0.05 0.06 0.063 0.083 0.067 0.047
Edge 2.5 0.45 2.07 0.99 0.11 0.36 0.31 0.36 0.33 0.29 0.26 0.33 0.35 0.31
Path 5 4.92 14.80 1.03 0.34 4.32 4.53 4.55 4.76 3.10 0.95 0.38 3.87 3.08
DBLP Node 0.2 5.56 418.07 18.70 0.55 6.98 7.98 8.22 9.67 1.66 31.32 10.48 5.73 3.03
Edge 0.2 8.76 414.47 22.57 0.90 10.07 14.76 12.89 12.32 3.53 33.91 14.46 8.87 5.27
Path 0.2 5.44 236.82 19.61 0.71 6.87 8.24 8.39 8.13 2.17 31.13 9.37 5.33 3.01
Yelp Node 0.1 2.56 6.10 13.84 1.02 6.81 1.19 0.96 1.16 1.12 1.84 1.68 1.02 6.58
Edge 1 19.42 67.55 16.88 2.01 13.61 8.58 8.97 9.17 6.93 30.23 10.22 10.27 6.03
Path 1 56.97 130.3 14.35 1.53 23.48 15.97 16.61 19.03 7.35 37.81 9.95 25.61 19.94